# Daya Bay Collaboration

### v1

@@ Abstract
In this Letter, the first measurement of the oscillation amplitude and frequency of reactor $\overline{\nu}_{e}$ at km-scale baselines using only the $n$H channel is reported. The analysis is performed with 1958 days of data collected by Daya Bay from 24 December 2012 to 30 August 2017, resulting in an $n$H sample that is $3.1$ times larger than the one used previously [10]. Two analyses relying on different calibration approaches, selections, background subtraction techniques, assessment of systematic uncertainties, and fitting methods are employed that yield consistent results. The large statistics, combined with improvements in the understanding of the systematic uncertainties, provides a $\sin^{2}2\theta_{13}$ precision surpassed only by Daya Bay’s $n$Gd result [7]. An $n$H-only measurement of the mass-squared difference $\Delta m^{2}_{32}$ is reported for the first time.
@@
Measuring the oscillation parameters involves performing a relative comparison of the rate and energy spectrum of reactor $\overline{\nu}_{e}$ at multiple detectors. This requires correcting for instabilities and differences in energy response within each detector (spatial non-uniformities), and establishing a common energy scale for all ADs. The two analyses used in this publication, referred to as A and B, begin from the reconstruction scheme described in Ref. [14], which relies primarily on spallation neutron (SPN) events tagged by $n$Gd capture to calibrate the energy response and is optimized for the GdLS region. To improve the performance, most importantly in the LS region, Analysis A calibrates the non-uniform energy response using the $2.2$ MeV signal from $n$H-SPN events. This analysis uses the fitted 2.2 MeV peak of $n$H-SPN events to align the energy scale of the ADs. The fit is done with a new electromagnetic calorimeter function that decouples energy leakage and detector non-uniformity effects [15]. Analysis B uses the $\alpha$ energy peak from sequential $\beta$-$\alpha$ decays of ${}^{214}\rm{Bi}$-${}^{214}\rm{Po}$ from naturally occurring 238U and 232Th to map the non-uniform response. The response is normalized relative to the central GdLS volume so as to rely on the same energy scale calibrations used in Daya Bay’s $n$Gd analysis [7]. After calibration, the energy scale difference among ADs is less than 0.30% (0.35%) for Analysis A (B), and the residual energy response non-uniformity within the ADs is less than $0.5\%$ for each analysis.
@@
Table 2 summarizes all AD-uncorrelated uncertainties for both analyses. The uncertainty in the number of protons for the target volumes is the same as in Ref. [10]. The uncertainty introduced by the minimum prompt-delayed time separation cut of $1~{}\mu\mathrm{s}$ is re-evaluated as in Ref. [10] and found to remain at $0.10\%$, while the $1500$ $\mu\mathrm{s}$ upper limit has a negligible inefficiency and thus no associated uncertainty.
@@
Table 2: AD-uncorrelated systematic uncertainties. The row of ‘Target protons’ accounts for relative differences in target mass whereas the next four account for differences in efficiency of the various selection efficiencies.
@@
The oscillation parameters are extracted from a $\chi^{2}$ fit that takes into account the observed $\overline{\nu}_{e}$ rate and spectrum of each AD together with the survival probability of Eq. (1). Analysis A performs a simultaneous fit of the data in all ADs allowing for correlated variations in the rate and spectrum of the reactor $\overline{\nu}_{e}$ prediction to account for any mismodeling. Two independent fitting programs A.1 and A.2 are constructed. Analysis B follows a strategy similar to Ref. [25] where a prediction at the far hall is made directly from the observation at the near halls, minimizing the dependence on reactor $\overline{\nu}_{e}$ prediction models, confirming the results from Analysis A. In both analyses, nuisance parameters are introduced in the $\chi^{2}$ definition to account for reactor-related, detector-related, and background-related uncertainties. The best-fit values of both analyses are reported in Table 3, where the fitting results of A.1 and A.2 are almost identical and are thus not repeated. The results of A and B are consistent when taking the analysis strategy difference into account. There is also good consistency with other estimates, most notably the latest one from Daya Bay using the $n$Gd sample [7].
@@
Figure 3 shows a comparison between the observed and expected prompt energy spectra in the far hall with no oscillation and with the best-fit oscillation parameters of Analysis A. Fig. 4 shows the best-fit result and the $68.3\%$, $95.5\%$, and $99.7\%$ confidence level allowed regions in the $\Delta m^{2}_{32}$ vs. $\sin^{2}2\theta_{13}$ plane. The statistical uncertainty of $\sin^{2}2\theta_{13}$ and $\Delta m^{2}_{\mathrm{ee}}$ accounts for about $47\%$ and $64\%$ of their total uncertainties, respectively.
The significant difference in selection criteria and in signal and background features makes the systematic uncertainties of this $\sin^{2}2\theta_{13}$ measurement virtually independent from those of the $n$Gd study. Using the result from Analysis A, the combined, weighted-average value of $\sin^{2}2\theta_{13}$ yields $0.0833\pm 0.0022$, which represents an $8\%$ improvement in precision with respect to the $n$Gd estimate alone that was obtained with 3158 days of data [7]. The combination of $\Delta m^{2}_{32}$ results suffers from more correlations between the $n$Gd and $n$H analyses, requiring more careful treatment.

### v2

@@ Abstract
In this Letter, the first measurement of the oscillation amplitude and frequency of reactor $\overline{\nu}_{e}$ at km-scale baselines using only the $n$H channel is reported. The analysis is performed with 1958 days of data collected by Daya Bay from 24 December 2012 to 30 August 2017, resulting in an $n$H sample that is $3.1$ times of the previous one [10]. Two analyses relying on different calibration approaches, selections, background subtraction techniques, assessment of systematic uncertainties, and fitting methods are employed that yield consistent results. The large statistics, combined with improvements in the understanding of the systematic uncertainties, provides a $\sin^{2}2\theta_{13}$ precision surpassed only by Daya Bay’s $n$Gd result [7]. An $n$H-only measurement of the mass-squared difference $\Delta m^{2}_{32}$ is reported for the first time.
@@
Measuring the oscillation parameters involves performing a relative comparison of the rate and energy spectrum of reactor $\overline{\nu}_{e}$ at multiple detectors. This requires correcting for instabilities and differences in energy response within each detector (spatial non-uniformities  [^1] The correction takes into account spatial non-uniformities due primarily to time-dependent changes in the number and location of inoperative channels in individual ADs.), and establishing a common energy scale for all ADs. The two analyses used in this publication, referred to as A and B, begin from the reconstruction scheme described in Ref. [15], which relies primarily on spallation neutron (SPN) events tagged by $n$Gd capture to calibrate the energy response and is optimized for the GdLS region. To improve the performance, most importantly in the LS region, Analysis A calibrates the non-uniform energy response using the $2.2$ MeV signal from $n$H-SPN events. This analysis uses the fitted 2.2 MeV peak of $n$H-SPN events to align the energy scale of the ADs. The fit is done with a new electromagnetic calorimeter function that decouples energy leakage and detector non-uniformity effects [16]. Analysis B uses the $\alpha$ energy peak from sequential $\beta$-$\alpha$ decays of ${}^{214}\rm{Bi}$-${}^{214}\rm{Po}$ from naturally occurring 238U and 232Th to map the non-uniform response. The response is normalized relative to the central GdLS volume so as to rely on the same energy scale calibrations used in Daya Bay’s $n$Gd analysis [7]. After calibration, the energy scale difference among ADs is less than 0.30% (0.35%) for Analysis A (B), and the residual energy response non-uniformity within the ADs is less than $0.5\%$ for each analysis.
@@
Table 2 summarizes all AD-uncorrelated uncertainties for both analyses. The total masses of GdLS, LS, IAV, OAV, and mineral oil were measured during the assembly of each AD, and the hydrogen mass fractions were also determined for each chemical component. The uncertainties in these measurements are propagated to the uncertainty in the efficiency-weighted number of target protons [10]. The uncertainty introduced by the minimum prompt-delayed time separation cut of $1~{}\mu\mathrm{s}$ is re-evaluated as in Ref. [10] and found to remain at $0.10\%$, while the $1500$ $\mu\mathrm{s}$ upper limit has a negligible inefficiency and thus no associated uncertainty.
@@
Table 2: AD-uncorrelated systematic uncertainties. The row of ‘Target protons’ accounts for relative differences in target mass whereas the next four account for differences in various selection efficiencies among ADs.
@@
The oscillation parameters are extracted from a $\chi^{2}$ fit that takes into account the observed $\overline{\nu}_{e}$ rate and spectrum of each AD together with the survival probability of Eq. (1). Analysis A performs a simultaneous fit of the data in all ADs allowing for correlated variations in the rate and spectrum of the reactor $\overline{\nu}_{e}$ prediction to account for any mismodeling. Two independent fitting programs A.1 and A.2 are constructed. Analysis B follows a strategy similar to Ref. [26] where a prediction at the far hall is made directly from the observation at the near halls, minimizing the dependence on reactor $\overline{\nu}_{e}$ prediction models, confirming the results from Analysis A. In both analyses, nuisance parameters are introduced in the $\chi^{2}$ definition to account for reactor-related, detector-related, and background-related uncertainties. The best-fit values of both analyses are reported in Table 3, where the fitting results of A.1 and A.2 are almost identical and are thus not repeated. The results of Analyses A and B are consistent. There is also good consistency with other estimates, most notably the latest one from Daya Bay using the $n$Gd sample [7].
@@
The result of Analysis A which achieves a better precision is chosen as nominal. Figure 3 shows a comparison between the observed and expected prompt energy spectra in the far hall with no oscillation and with the best-fit oscillation parameters of Analysis A. Fig. 4 shows the best-fit result and the $68.3\%$, $95.5\%$, and $99.7\%$ confidence level allowed regions in the $\Delta m^{2}_{32}$ vs. $\sin^{2}2\theta_{13}$ plane. The statistical uncertainty of $\sin^{2}2\theta_{13}$ and $\Delta m^{2}_{\mathrm{ee}}$ accounts for about $47\%$ and $64\%$ of their total uncertainties, respectively.
The significant difference in selection criteria and in signal and background features makes the systematic uncertainties of this $\sin^{2}2\theta_{13}$ measurement virtually independent from those of the $n$Gd study. Using the result from Analysis A, the combined, weighted-average value of $\sin^{2}2\theta_{13}$ yields $0.0833\pm 0.0022$, which represents an $8\%$ improvement in precision with respect to the $n$Gd estimate alone that was obtained with 3158 days of data [7]. The combination of $\Delta m^{2}_{32}$ results suffers from more correlations between the $n$Gd and $n$H analyses, requiring additional study.